INVESTIGATION OF THE ROVIBRATIONAL ENERGY PATTERN IN ACETALDEHYDE IN THE MID AND NEAR INFRARED RANGES

\underline{I. KLEINER}, Laboratoire de Photophysique Moléculaire, Batiment 210, Université PARIS 11, Campus d’ORSAY, 91405 ORSAY CEDEX, FRANCE; M. HEPP, R. GEORGES, and M. HERMAN, Laboratoire de Chimie Physique Moléculaire, CP160/09, Université Libre de Bruxelles, B-1050, Belgium.

Three ranges of the infrared and near infrared absorption spectrum of CH_{3}CHO have been recorded using a Fourier transform interferometer, under various experimental conditions: between 1300 and 3100 cm^{−1} at a resolution of 0.01 cm^{−1} under jet cooled conditions, between 1400 and 3200 cm^{−1} at a resolution of 0.008 cm^{−1} at room temperature, and between 4000 and 8000 cm^{−1} at a resolution of 0.03 cm^{−1} at room temperature. We are investigating specific bands in those ranges, all presenting problems of different nature, severely limiting the analysis in each case. We are studying ν_{5} and ν_{12} respectively located at 1429.91 and 1435.83 cm^{−1}. They are two methyl bending fundamentals, with strong Coriolis interaction^{a}. We are also focusing on the aldehydic CH stretch fundamental band (ν_{3}), located around 2716 cm^{−1} and known to be in Fermi-type interaction with 2ν_{6}^{b}. Eventually, we are considering the two first overtones of the aldehydic CH stretch, observed around 5320 and 7810 cm^{−1}, presenting a dense structure.

aH. Hollenstein, Mol. Phys., 39 (1980) 1013.

bH. Hollenstein and Hs. H. Günthard, Spectrochim. Acta 27A, 2027 1971.